# On Secret-Message Transmission by Echoing Encrypted Probes

### v1

@@ III-B Phase 2 of G-STEEP
$\mathbf{R}_{\Delta\mathbf{p}^{\prime}}\doteq\mathbb{E}\{\Delta\mathbf{p}^{\prime}\Delta\mathbf{p}^{{}^{\prime}H}\}$
@@
III-C Effective channel capacity from Bob to Alice
The effective channel capacity from Bob to Alice relative to $\mathbf{s}$ (in bits per round-trip symbol interval) is therefore
@@ III-C Effective channel capacity from Bob to Alice
where $N_{A|B}$ and $D_{A|B}$ are defined in the obvious way. Here $|A|$ denotes the determinant of $A$. Notice that $\mathbf{s},\mathbf{x}_{A},\mathbf{y}_{A}$ are jointly Gaussian so that the 2nd and 3rd equalities in (III-C) hold. Note that for $p_{A}\to 0\texttt{ or }\infty$, $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}\to\mathbf{0}$ and hence $C_{A|B,G}\to\log|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}+\mathbf{I}_{n_{B}}|$.
III-D Effective channel capacity from Bob to Eve
@@
where $\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}\doteq\mathbb{E}\{\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\}$. Let $\mathbf{Q}\doteq[\mathbf{V}_{BA},\mathbf{V}_{BA}^{\perp}]\in\mathcal{C}^{n_{A}\times n_{A}}$ be a unitary matrix. Then
@@
Again we have applied the jointly Gaussian nature of $\mathbf{s},\mathbf{y}_{EA},\mathbf{y}_{EB}$ for the 2nd and 3rd equalities in (III-D).
III-E Secrecy capacity of G-STEEP
@@
An achievable secrecy rate of G-STEEP based on the effective wiretap-channel system from Bob to Alice against Eve (in bits per round-trip symbol interval) is
$C_{s,G}\doteq(I(\mathbf{s};\{\mathbf{x}_{A},\mathbf{y}_{A}\})-I(\mathbf{s};\{\mathbf{y}_{EA},\mathbf{y}_{EB}\}))^{+}$
@@
where $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}$ is given in (III-B), and $\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}$ is given in (40).
@@
This follows from the WTC theory for Gaussian channels [2] with respect to the message $\mathbf{s}$ from Bob, and the previous results shown in (III-C) and (III-D). ∎
@@
One may argue that the secret-key capacity based on $\{\mathbf{x}_{A},\mathbf{y}_{A}\}$ at Alice, $\{\mathbf{y}_{B},\mathbf{s}\}$ at Bob and $\{\mathbf{y}_{EA},\mathbf{y}_{EB}\}$ at Eve after both phases of G-STEEP (and using additional and iterative operations for information reconciliation and privacy amplification via public communications) should be larger than or equal to $C_{s,G}$. But this paper will not pursue that direction of exploration because WTC transmission is relatively simple and its secrecy rate over the effective WTC system induced by STEEP is already promising as discussed in the rest of this section.
@@
If $n_{B}=1$ and $\mathbf{H}_{AB}$, $\mathbf{H}_{BA}$ and $\mathbf{H}_{EB}$ are replaced by $\mathbf{h}_{AB}$, $\mathbf{h}_{BA}$ and $\mathbf{h}_{EB}$ (similarly for their scaled versions), then $C_{s,G}\doteq\left(C_{A|B,G}-C_{E|B,G}\right)^{+}$ with
@@
This follows from Theorem 1. In particular, $\mathbf{T}$ in (III-D) is now reduced to the scalar $t$. ∎
We see that for the case of $n_{B}=1$, the effects of $\mathbf{h}_{AB}^{\prime}$, $\mathbf{h}_{BA}^{\prime}$ and $\mathbf{h}_{EB}^{\prime}$ on $C_{s,G}$ are only through their norms. The effect of $\mathbf{H}_{EA}^{\prime}$ on $C_{s,G}$ is only through the scalar $t$.
III-F Properties of $C_{s,G}$
Lemma 1
Proof:
This follows from Theorem 1 in [11] where Maurer’s lower and upper bounds, generalized in [2] and further applied asymptotically to continuous sources via generalized mutual information, are used. (Some earlier attempt such as [26] to improve Maurer’s lower bound is not necessary here.) $C_{key}$ here is $\xi_{B}$ in equation (8) in [11] with constant channel matrices, and $\mathbf{H}_{EA}$ and $\frac{p_{A}}{n_{A}}$ here are $\mathbf{G}_{A}$ and $\gamma_{BA}$ in [11]. Furthermore, $\lambda_{B}$ and $\lambda_{EA}$ in [11] are both normalized to be one here. ∎
Note that $C_{key}$ is the maximum secret-key rate achievable based on the data sets generated in phase 1 of G-STEEP at Alice, Bob and Eve and through communications in the public network (where Eve has access to all communications). So, if Eve’s receive channel from Bob is no weaker than Alice’s receive channel from Bob (in phase 2 of G-STEEP), we should expect $C_{s,G}\leq C_{key}$. If $C_{s,G}$ approaches $C_{key}$ under high powers, we can say that $C_{s,G}$ is optimal against strong Eve under high powers.
But if Eve’s receive channel from Bob is weaker than that at Alice, it is possible to have $C_{s,G}>C_{key}$. But one should not be excited by this situation. We know that $C_{key}$ is based on the assumption that all communications for secret key generation are done in the public domain. If any of these communications are not public, the resulting $C_{key}$ would be higher. So, for a meaningful comparison between $C_{s,G}$ and $C_{key}$, we should assume that Eve’s receive channel in phase 2 of G-STEEP is no weaker than that at Alice.
@@ III-F Properties of $C_{s,G}$
$\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{s,G}=\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{key}$
$=\min(n_{B},(n_{A}-n_{E})^{+}),$ (48)
i.e., $\texttt{DoF}(C_{s,G})=\texttt{DoF}(C_{key})$. Namely, $C_{s,G}$ is optimal in DoF.
Proof:
See Appendix-A. ∎
Proposition 2
$\lim_{p_{A}\to\infty}(\lim_{p_{B}\to\infty}C_{s,G})=\lim_{p_{A}\to\infty}C_{key}$
Namely, $C_{s,G}$ is optimal (against strong Eve) asymptotically as $p_{A}\to\infty$ and $\frac{p_{B}}{p_{A}}\to\infty$.
Proof:
See Appendix-B. ∎
The above proposition is also intuitively justified if we think of $\frac{p_{B}}{p_{A}}\to\infty$ as somewhat similar to the case where phase 2 of G-STEEP only uses public communications and also think of $p_{A}\to\infty$ as somewhat similar to the case where the encryption in phase 2 is done via a modulo sum between two discrete random variables. In other words, for $\frac{p_{B}}{p_{A}}\to\infty$, both Alice and Eve would receive the same $\sqrt{p_{B}}(\mathbf{\hat{p}}+\mathbf{s})$ from Bob, i.e., the phase 2 would be via a public channel. For $p_{B}\to\infty$, $\sqrt{p_{B}}(\mathbf{\hat{p}}+\mathbf{s})=\sqrt{p_{B}}\mathbf{\hat{p}}+\sqrt{p_{B}}\mathbf{s}$ is a sum between $\sqrt{p_{B}}\mathbf{\hat{p}}$ and $\sqrt{p_{B}}\mathbf{s}$ (which would be virtually uniformly distributed), and this sum would be like a modulo-sum with an infinite modulo. Then the encryption lemma would suggest that in the case of $p_{A}\to\infty$ and $\frac{p_{B}}{p_{A}}\to\infty$, $C_{key}=C_{s,G}$. Again, the above discussion is no proof but only an intuition to make an intuitive sense of the result actually proven in Appendix-B.
Since the limit in (2) is always positive (unless $\mathbf{H}_{EA}$ has an infinite norm), this proposition also suggests that for a sufficiently large (but finite) $p_{A}$ and a sufficiently large (but finite) $\frac{p_{B}}{p_{A}}$, $C_{s,G}$ is positive. We will see a more specific case of this next.
III-G The special case of $n_{A}=n_{B}=1$
$C_{s,G}=\left[\log\left(1+\frac{b/2}{bA_{1}/2+1}\right)-\log\left(1+\frac{\beta b/2}{\beta bA_{2}/2+1}\right)\right]^{+}$ (50)
with $a\doteq S_{BA}\doteq p_{A}|h_{BA}|^{2}=|h_{BA}^{\prime}|^{2}$ and $b\doteq S_{AB}\doteq p_{B}|h_{AB}|^{2}=|h_{AB}^{\prime}|^{2}$. Also $A_{1}=\frac{a}{(a+1)^{2}}$ and $A_{2}=A_{1}\frac{(a+\alpha a+1)}{\alpha a+1}$ with $\alpha\doteq\frac{S_{EA}}{S_{BA}}$ and $\beta\doteq\frac{S_{EB}}{S_{AB}}$. Furthermore, $S_{EA}\doteq p_{A}\|\mathbf{h}_{EA}\|^{2}=\|\mathbf{h}_{EA}^{\prime}\|^{2}$ and $S_{EB}\doteq p_{B}\|\mathbf{h}_{EB}\|^{2}=\|\mathbf{h}_{EB}^{\prime}\|^{2}$. Note that $A_{2}>A_{1}$ and they are invariant to $b$.
In particular, if Eve’s (raw) channel is stronger than users’ (raw) channel in phase 1, then $\alpha>1$; and if Eve’s (raw) channel is stronger than users’ (raw) channel in phase 2, then $\beta>1$. It is obvious that $C_{s,G}$ increases as $\alpha$ and/or $\beta$ decrease.
If $\alpha\geq 1$ and $\beta\geq 1$, all conventional WTC schemes either from Alice to Bob or from Bob to Alice have zero secrecy capacity. (In the MIMO case, a similar conclusion can be drawn if $n_{E}\geq n_{A}\geq n_{B}$ and the large-scale fading at Eve is smaller than those at Alice and Bob.)
Proposition 3
Assume $n_{A}=n_{B}=1$. Then $C_{s,G}>0$ if and only if
$b>\bar{b}\doteq\frac{2(\beta-1)}{\beta(A_{2}-A_{1})}$
$=\frac{2(\beta-1)(a+1)^{2}(\alpha a+1)}{\beta a^{2}}.$ (51)
Also for $\beta>1$, $C_{s,G}$ increases as $b$ increases.
@@ III-G The special case of $n_{A}=n_{B}=1$
$\bar{b}\approx 2\frac{\beta-1}{\beta}\alpha a=\mathcal{O}(a)$ (52)
Hence, for $\alpha>1$, $\beta>1$ and a given $b$ (such as 20dB to 30dB in practice), the optimal value of $a$ to maximize $C_{s,G}$ is generally in between zero and $b$.
In practice, we can utilize (3) to ensure a positive secrecy rate whenever an upper bound on $\alpha$ is available. In the case of random fading channels, the probability for (3) not to hold can be kept small by keeping a large ratio of $p_{B}$ over $p_{A}$ [29].
III-G1 Comparison to $C_{key}$

### v2

@@ III-C Analysis of the signal received by Alice in phase 2
$\mathbf{R}_{\Delta\mathbf{p}^{\prime}}\doteq\mathbb{E}\{\Delta\mathbf{p}^{\prime}\Delta\mathbf{p}^{\prime H}\}$
@@
III-D Effective channel capacity from Bob to Alice
Lemma 1
If $\mathbf{x}\in\mathbb{C}^{n_{x}}$ and $\mathbf{y}\in\mathbb{C}^{n_{y}}$ are joint (non-singular) circular complex Gaussian with zero means and covariance matrices $\mathbf{R}_{x}$ and $\mathbf{R}_{y}$ respectively, then $I(\mathbf{x};\mathbf{y})=I(\mathbf{x};\mathbf{\hat{x}})=\log\frac{|\mathbf{R}_{x}|}{|\mathbf{R}_{x|y}|}$, where $\mathbf{\hat{x}}$ is the MMSE estimate of $\mathbf{x}$ from $\mathbf{y}$, and $\mathbf{R}_{x|y}$ is the MSE matrix of $\mathbf{\hat{x}}$.
Proof:
Let $\mathbf{z}=[\mathbf{x}^{T},\mathbf{y}^{T}]^{T}$. Then $\mathbf{z}$ is Gaussian with the covariance matrix
$\mathbf{R}_{z}=\left[\begin{array}[]{cc}\mathbf{R}_{x}&\mathbf{R}_{xy}\\ \mathbf{R}_{xy}^{H}&\mathbf{R}_{y}\end{array}\right].$ (19)
It follows that the PDF of $\mathbf{x}$ given $\mathbf{y}$ is
$f(\mathbf{x}|\mathbf{y})=\frac{f(\mathbf{x},\mathbf{y})}{f(\mathbf{y})}=\frac{\frac{1}{\pi^{n_{x}+n_{y}}|\mathbf{R}_{z}|}\exp(-\mathbf{z}^{H}\mathbf{R}_{z}^{-1}\mathbf{z})}{\frac{1}{\pi^{n_{y}}|\mathbf{R}_{y}|}\exp(-\mathbf{y}^{H}\mathbf{R}_{y}^{-1}\mathbf{y})}$ (20)
Using the block matrix properties of $\mathbf{R}_{z}^{-1}$ and $|\mathbf{R}_{z}|$, one can verify that
$f(\mathbf{x}|\mathbf{y})=\frac{1}{\pi^{n_{x}}|\mathbf{R}_{x|y}|}\exp\left(-(\mathbf{x}-\mathbf{\hat{x}})^{H}\mathbf{R}_{x|y}^{-1}(\mathbf{x}-\mathbf{\hat{x}})\right)$ (21)
with $\mathbf{\hat{x}}\doteq\mathbf{R}_{xy}\mathbf{R}_{y}^{-1}\mathbf{y}$ and $\mathbf{R}_{x|y}\doteq\mathbf{R}_{x}-\mathbf{R}_{xy}\mathbf{R}_{y}^{-1}\mathbf{R}_{xy}^{H}$. We see that this $\mathbf{\hat{x}}$ is the MMSE estimate of $\mathbf{x}$ from $\mathbf{y}$ because $\mathbb{E}\{\mathbf{x}|\mathbf{y}\}=\mathbf{\hat{x}}$, and this $\mathbf{R}_{x|y}$ is the MSE matrix of $\mathbf{\hat{x}}$. Finally, we know $I(\mathbf{x};\mathbf{y})=h(\mathbf{x})-h(\mathbf{x}|\mathbf{y})=\log|\mathbf{R}_{x}|-\log|\mathbf{R}_{x|y}|$. A constant term in each of the differential entropies $h(\mathbf{x})$ and $h(\mathbf{x}|\mathbf{y})$ canceled each other. We also see that $\mathbf{\hat{x}}$ is a sufficient statistic of $\mathbf{y}$ for $\mathbf{x}$, and hence $I(\mathbf{x};\mathbf{y})=I(\mathbf{x};\mathbf{\hat{x}})$. ∎
The virtual channel from $\mathbf{s}$ to $\mathbf{\hat{s}}_{A}$ is called here the effective channel from Bob to Alice relative to $\mathbf{s}$, the capacity of which (in bits per round-trip symbol interval) is therefore
@@ III-D Effective channel capacity from Bob to Alice
where $N_{A|B}$ and $D_{A|B}$ are defined in the obvious way. Here $|A|$ denotes the determinant of $A$. Notice that $\mathbf{s},\mathbf{x}_{A},\mathbf{y}_{A}$ are jointly Gaussian so that the 2nd and 3rd equalities in (III-D) hold. Note that for $p_{A}\to 0\texttt{ or }\infty$, $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}\to\mathbf{0}$ and hence $C_{A|B,G}\to\log|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}+\mathbf{I}_{n_{B}}|$.
III-E Effective channel capacity from Bob to Eve
To determine the effective channel capacity from Bob to Eve, we need to determine the MSE matrix of the MMSE estimate of $\mathbf{s}$ based on all signals observed by Eve in phases 1 and 2.
@@ III-E Effective channel capacity from Bob to Eve
where $\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}\doteq\mathbb{E}\{\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\}$. Let $\mathbf{Q}\doteq[\mathbf{V}_{BA},\mathbf{V}_{BA}^{\perp}]\in\mathbb{C}^{n_{A}\times n_{A}}$ be a unitary matrix. Then
@@
Again we have applied the jointly Gaussian nature of $\mathbf{s},\mathbf{y}_{EA},\mathbf{y}_{EB}$ for the 2nd and 3rd equalities in (III-E).
III-F Secrecy rate of G-STEEP
@@
An achievable secrecy rate of G-STEEP based on the effective wiretap-channel system from Bob to Alice against Eve (in bits per round-trip symbol interval or two complex channel uses) is
$R_{s,G}\doteq(I(\mathbf{s};\{\mathbf{x}_{A},\mathbf{y}_{A}\})-I(\mathbf{s};\{\mathbf{y}_{EA},\mathbf{y}_{EB}\}))^{+}$
@@
where $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}$ is given in (III-C), and $\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}$ is given in (47).
@@
This follows from the WTC theory for Gaussian signaling over Gaussian noise channels [2] with respect to the message vector $\mathbf{s}$ from Bob, and the previous results shown in (III-D) and (III-E). ∎
III-G Properties of secrecy rate of G-STEEP
Lemma 2
Proof:
This lemma is a special case of Theorem 1 in [11] where Maurer’s lower and upper bounds, generalized in [2] and further applied asymptotically to continuous sources via generalized mutual information, are used. (Some earlier attempt such as [26] to improve Maurer’s lower bound is not necessary here.) Specifically, the lower bound $I(\mathbf{x}_{A};\mathbf{y}_{B})-I(\mathbf{y}_{B};\mathbf{y}_{E,A})$ of $C_{key}$, which reduces to $[h(\mathbf{y}_{B})-h(\mathbf{y}_{B}|\mathbf{x}_{A})]-[h(\mathbf{y}_{B})-h(\mathbf{y}_{B}|\mathbf{y}_{E,A})]=h(\mathbf{y}_{B}|\mathbf{y}_{E,A})-h(\mathbf{y}_{B}|\mathbf{x}_{A})$, equals the upper bound $I(\mathbf{x}_{A};\mathbf{y}_{B}|\mathbf{y}_{E,A})=h(\mathbf{y}_{B}|\mathbf{y}_{E,A})-h(\mathbf{y}_{B}|\mathbf{x}_{A},\mathbf{y}_{E,A})=h(\mathbf{y}_{B}|\mathbf{y}_{E,A})-h(\mathbf{y}_{B}|\mathbf{x}_{A})$ where the last equation follows from the fact that $\mathbf{y}_{B}$ and $\mathbf{y}_{E,A}$ are independent of each other when conditioned on $\mathbf{x}_{A}$. One can verify that (51) follows from $C_{key}=h(\mathbf{y}_{B}|\mathbf{y}_{E,A})-h(\mathbf{y}_{B}|\mathbf{x}_{A})$, which is also $\xi_{B}$ in equation (8) in [11] with constant channel matrices, and $\mathbf{H}_{EA}$ and $\frac{p_{A}}{n_{A}}$ here are $\mathbf{G}_{A}$ and $\gamma_{BA}$ in [11]. Furthermore, $\lambda_{B}$ and $\lambda_{EA}$ in [11] are both normalized to be one here. Alternatively, see [33]. ∎
Note that $C_{key}$ is the maximum secret-key rate achievable based on the data sets (among all possible statistical distributions of the probes) generated in phase 1 of G-STEEP at Alice, Bob and Eve and through communications in the public network to which Eve has full access. So, if Eve’s receive channel from Bob is no weaker than Alice’s receive channel from Bob (in phase 2 of G-STEEP), we should expect $R_{s,G}\leq C_{key}$. A precise statement is shown later in Proposition 4. If $R_{s,G}$ approaches $C_{key}$ under high powers, we can say that $R_{s,G}$ is optimal against strong Eve under high powers.
But if Eve’s receive channel from Bob is weaker than that at Alice, it is possible to have $R_{s,G}>C_{key}$. But one should not be excited by this situation. We know that $C_{key}$ is based on the assumption that all communications for secret key generation are done in the public domain. If any of these communications are not public (i.e., secure or partially secure), the resulting $C_{key}$ would be higher. So, for a meaningful comparison between $R_{s,G}$ and $C_{key}$, we should assume that Eve’s receive channel in phase 2 of G-STEEP is no weaker than that at Alice.
@@ III-G Properties of secrecy rate of G-STEEP
One may argue that the secret-key capacity $C_{key}^{\prime}$ based on $\{\mathbf{x}_{A},\mathbf{y}_{A}\}$ at Alice, $\{\mathbf{y}_{B},\mathbf{s}\}$ at Bob and $\{\mathbf{y}_{EA},\mathbf{y}_{EB}\}$ at Eve after both phases of G-STEEP (and using additional and iterative operations for information reconciliation and privacy amplification via public communications) should always be larger than or equal to $R_{s,G}$. But the analysis of $C_{key}^{\prime}$ is more involved, and there is a gap between its lower and upper bounds.
Proposition 1
$\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}R_{s,G}=\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{key}$
$=\min(n_{B},(n_{A}-n_{E})^{+}),$ (52)
i.e., $\texttt{DoF}(R_{s,G})=\texttt{DoF}(C_{key})$. Namely, $R_{s,G}$ is optimal in DoF.
Proof:
See Appendix-A. ∎
Proposition 2
$\lim_{p_{A}\to\infty}(\lim_{p_{B}\to\infty}R_{s,G})=\lim_{p_{A}\to\infty}C_{key}$
Namely, $R_{s,G}$ is optimal (against strong Eve) asymptotically as $p_{A}\to\infty$ and $\frac{p_{B}}{p_{A}}\to\infty$.
Proof:
See Appendix-B. ∎
The above proposition is also intuitively justified if we think of $\frac{p_{B}}{p_{A}}\to\infty$ as somewhat similar to the case where phase 2 of G-STEEP only uses public communications and also think of $p_{A}\to\infty$ as somewhat similar to the case where the encryption in phase 2 is done via a modulo sum between two discrete random variables. In other words, for $\frac{p_{B}}{p_{A}}\to\infty$, both Alice and Eve would receive the same $\sqrt{p_{B}}(\mathbf{\hat{p}}+\mathbf{s})$ from Bob, i.e., the phase 2 would be via a public channel. For $p_{B}\to\infty$, $\sqrt{p_{B}}(\mathbf{\hat{p}}+\mathbf{s})=\sqrt{p_{B}}\mathbf{\hat{p}}+\sqrt{p_{B}}\mathbf{s}$ is a sum between $\sqrt{p_{B}}\mathbf{\hat{p}}$ and $\sqrt{p_{B}}\mathbf{s}$ (which would be virtually uniformly distributed), and this sum would be like a modulo-sum with an infinite modulo. Then the encryption lemma would suggest that in the case of $p_{A}\to\infty$ and $\frac{p_{B}}{p_{A}}\to\infty$, $C_{key}=R_{s,G}$. Again, the above discussion is no proof but only an intuition to make an intuitive sense of the result actually proven in Appendix-B.
Since the limit in (2) is always positive (unless $\mathbf{H}_{EA}$ has an infinite norm), this proposition also suggests that for a sufficiently large (but finite) $p_{A}$ and a sufficiently large (but finite) $\frac{p_{B}}{p_{A}}$, $R_{s,G}$ is positive. We will see a more specific case of this next.
@@
If $n_{B}=1$ and $\mathbf{H}_{AB}$, $\mathbf{H}_{BA}$ and $\mathbf{H}_{EB}$ are replaced by $\mathbf{h}_{AB}$, $\mathbf{h}_{BA}$ and $\mathbf{h}_{EB}$ (similarly for their scaled versions), then $R_{s,G}\doteq\left(C_{A|B,G}-C_{E|B,G}\right)^{+}$ with
@@
This follows from Theorem 1. In particular, $\mathbf{T}$ in (III-E) is now reduced to the scalar $t$. ∎
We see that for the case of $n_{B}=1$, the effects of $\mathbf{h}_{AB}^{\prime}$, $\mathbf{h}_{BA}^{\prime}$ and $\mathbf{h}_{EB}^{\prime}$ on $R_{s,G}$ are only through their norms. The effect of $\mathbf{H}_{EA}^{\prime}$ on $R_{s,G}$ is only through the scalar $t$.
III-H The special case of $n_{A}=n_{B}=1$
III-H1 Analysis of $R_{s,G}$
$R_{s,G}=\left[\log\left(1+\frac{b/2}{bA_{1}/2+1}\right)-\log\left(1+\frac{\beta b/2}{\beta bA_{2}/2+1}\right)\right]^{+}$ (57)
with $a\doteq S_{BA}\doteq p_{A}|h_{BA}|^{2}=|h_{BA}^{\prime}|^{2}$ and $b\doteq S_{AB}\doteq p_{B}|h_{AB}|^{2}=|h_{AB}^{\prime}|^{2}$. Also $A_{1}=\frac{a}{(a+1)^{2}}$ and $A_{2}=A_{1}\frac{(a+\alpha a+1)}{\alpha a+1}$ with $\alpha\doteq\frac{S_{EA}}{S_{BA}}$ and $\beta\doteq\frac{S_{EB}}{S_{AB}}$. Furthermore, $S_{EA}\doteq p_{A}\|\mathbf{h}_{EA}\|^{2}=\|\mathbf{h}_{EA}^{\prime}\|^{2}$ and $S_{EB}\doteq p_{B}\|\mathbf{h}_{EB}\|^{2}=\|\mathbf{h}_{EB}^{\prime}\|^{2}$. Note that $A_{1}<A_{2}<1$ and they are invariant to $b$.
In particular, if Eve’s (raw) channel is stronger than users’ (raw) channel in phase 1, then $\alpha>1$; and if Eve’s (raw) channel is stronger than users’ (raw) channel in phase 2, then $\beta>1$.
If $\alpha\geq 1$ and $\beta\geq 1$, all conventional WTC schemes either from Alice to Bob or from Bob to Alice have zero secrecy capacity.
@@ III-H1 Analysis of $R_{s,G}$
For $n_{A}=n_{B}=1$, $R_{s,G}>0$ if and only if
$b>\bar{b}\doteq\frac{2(\beta-1)}{\beta(A_{2}-A_{1})}=\frac{2(\beta-1)(a+1)^{2}(\alpha a+1)}{\beta a^{2}}.$ (58)
@@
$\bar{b}\approx 2\frac{\beta-1}{\beta}\alpha a=\mathcal{O}(\alpha a)$ (59)
In practice, one can utilize (58) to ensure a positive secrecy rate whenever an upper bound on $\alpha$ (not necessarily on $\beta$) is available. In the case of random fading channels, the probability for (58) not to hold can be kept small by keeping a large ratio of $p_{B}$ over $p_{A}$ [29].
One can also verify that $R_{s,G}$ increases as $\alpha$ and/or $\beta$ decrease; for $\beta>1$, $R_{s,G}$ increases as $b$ increases, but $R_{s,G}$ saturates as $b$ becomes large; and $R_{s,G}$ versus $a$ is not monotonic in general. For a given $b$, $R_{s,G}$ generally peaks at a value of $a$ in between zero and $b$.
III-H2 Comparison to $C_{key}$